OpenStudy (anonymous):

find the area of r that is bounded by the graph of y=x^2-1 and the graph of x=y^2

OpenStudy (anonymous):

is this for single or multi-variable calculus?

OpenStudy (amistre64):

do I dare? :)

OpenStudy (nowhereman):

It's single variable calculus.

OpenStudy (amistre64):

this would be easier if we do it with respect to the y axis right?

OpenStudy (anonymous):

I don't think so.

OpenStudy (amistre64):

I get pretty little intervals on the up and down tho ....

OpenStudy (anonymous):

Err. Oh. Yes, doing it with respect to y is best. I misread the equation.

OpenStudy (anonymous):

if this is for multi-variable, I would integrate w.r.t. x first, if it is for single- then use the little formula for area between two curves wr.t .y

OpenStudy (anonymous):

It's single variable.

OpenStudy (anonymous):

@ polpak, Is this your question?

OpenStudy (amistre64):

is this equivalent to: y=x^2 y = sqrt(x-1) ?? I think I did it right, but aint sure...

OpenStudy (amistre64):

y=sqrt(x+1) wrote down wrong.... im good

OpenStudy (anonymous):

\[f_1(y) = \sqrt{1+y}\] \[f_2(y) = y^2\]

OpenStudy (anonymous):

agreed

OpenStudy (amistre64):

I see the points on the paper, I just cant get to them :)

OpenStudy (amistre64):

x^4 -2x^3 -x +1 = 0 .... is this even right yet?

OpenStudy (amistre64):

can we modify this with trig? use polar coordinates? just wondering.. If we add the functions together, can we get the area that way?

OpenStudy (amistre64):

not adding i guess... maybe subtracting :)